OpenStudy (anonymous):

the letters r and represent polar coordinates. Write each equation using rectangular coordinates (x,y): r=cosƟ

OpenStudy (anonymous):

multiply both sides by r : r^2 = rcos(theta) x^2 +y^2 = x x^2 -x +1/4 +y^2 = 1/4 (add both sides by 1/4 to make it quadratic equation) (x-1/2)^2 +y^2 = 1/4 this is circle center (1/2,0) radius 1/2